OpenStudy (anonymous):

solve for x: 2/5x = 16

OpenStudy (anonymous):

solve for x..... so to leave x alone since the 5 is dividing pass it to the other side by multiplying. so, 2x= 16(5) which equals 2x=80... now since the 2 is multiplying the x just pass it to the other side dividing x=80/2, now solve so, x=40!!

OpenStudy (anonymous):

thankyou :)

OpenStudy (anonymous):

Is this 2/(5x) or (2/5) * x? I assume the x is in the denominator, with the 5. In which case, you multiply the 5x by both sides to get 80*5x = 2, then solve for x. That will give you a slightly different answer than the one above.

OpenStudy (anonymous):

the first one

OpenStudy (anonymous):

Sorry, I meant to write 16*5x = 80x, then you have 80x = 2, divide the 80 and you get 1/40. So, x = 1/40.